# SPACeR: Self-Play Anchoring with Centralized Reference Models

### v1

@@ 4.3 SPACeR Agents for Fast & Accurate Closed-Loop Planner Evaluation
A key application of humanlike agents is the closed-loop evaluation of ego-vehicle motion planners. In this setup, the planner under test controls the ego vehicle, while surrounding agents are driven by learned policies. This enables realistic modeling of inter-agent interactions, reactive behaviors, and humanlike decision-making—without risking safety.
We evaluate our SPACeR agents along two dimensions: realism and throughput. Realism is crucial for faithfully assessing planner performance, while high throughput enables large-scale evaluation, covering rare but critical corner cases.
To assess realism, we compare against two baselines: (i) open-loop log-replay evaluation, and (ii) closed-loop simulation using Cat-K Zhang et al. (2025b), a state-of-the-art traffic simulator on WOSAC. We evaluate a diverse set of planners: 22 self-play–trained policies, 10 sampling-based Frenet planners, and 10 IDM-based planners Treiber et al. (2000) (more details can be found in A.5). For each planner, we compute PDM scores Dauner et al. (2024) across multiple scenes under three traffic-simulation strategies: ground-truth logs (GT), Cat-K rollouts, and SPACeR agent policies. We then measure the correlation of PDM scores between strategies. As shown in fig. 2, our approach yields consistently lower correlations with GT replays when compared to Cat-K. We interpret this as evidence that our simulation is more reactive, and suppresses unrealistic planner behavior more effectively—particularly in collision scenarios. Qualitative examples (see fig. 2 and supplementary videos) further highlight that SPACeR agents respond naturally in diverse scenarios, avoiding collisions when reasonable and minimizing unrealistic off-road behaviors.
@@
In summary, SPACeR agents deliver both more realistic and significantly faster closed-loop planner evaluation, enabling reliable and scalable benchmarking of ego-motion planning policies.
@@ 5 Discussion
Simulating VRUs. Extending our framework to vulnerable road users (VRUs) such as pedestrians and cyclists remains challenging. Current simulators provide only limited infrastructure support, where VRUs are often penalized solely through off-road rewards without richer semantic context such as traffic lights or crosswalks. This misalignment makes it difficult to design meaningful rewards or evaluate VRU behavior without explicit goal conditioning. Addressing this gap will be important for more comprehensive multi-agent simulation.
@@ 6 Conclusion
Acknowledgments
The authors would like to thank Shubh Gupta for insightful discussions on the experimental design in this paper.
@@ Appendix A Appendix
A.1 Training hyperpareamters and implementation details of Self-play
The core PPO hyperparameters are summarized in table 3. When training without a reference model, we employ 600 parallel worlds on a single A100 (80 GB, PCIe) GPU. For human-regularized PPO and reference-model training, we reduce the number of worlds to 300, which leads to an approximately $2\times$ slowdown due to the lack of multi-GPU support in GPUDRIVE with madrona backend. With multi-GPU support, we would expect comparable throughput to the reference-free setting. Our policy/value network is configured with an input embedding dimension of 64, a hidden dimension of 128, and a dropout rate of 0.01. To reduce GPU memory consumption in the reference-model setting, we highlight two adjustments: (1) we cap the maximum number of unique scenarios per batch at 200 to increase training speed, and (2) we limit the maximum number of map elements per agent from 200 Kazemkhani et al. (2024) to 120. table 3
Table 3: PPO Training Hyperparameters.
@@ A.1 Training hyperpareamters and implementation details of Self-play
We follow the setup of human-regularized PPO Cornelisse & Vinitsky (2024), but adapt it to our tokenized trajectory action space instead of low-level control actions. To train the behavior cloning (BC) reference policy, we use observation–action pairs from the full Waymo Open Motion Dataset (WOMD) rather than the 200-scenario subset originally used, ensuring a stronger expert baseline. Our BC model is parameterized with roughly $2\times$ the capacity of the self-play policy network to increase expressiveness, achieving a validation accuracy of 92%. We implement BC training using the imitation package Gleave et al. (2022) and train for 60 epochs on the full dataset.
During HR-PPO training, we regularize the learned policy against the BC reference policy using a KL-divergence penalty with weight $\beta=0.01$. Larger values of $\beta$ destabilize training, while using toeknzied model can generally increase to $\beta=1.0$ w/ similar performance. In addition, when applying human-regularized PPO in the original action space, we found it necessary to clamp the minimum reference log-probability to $10^{-20}$ to avoid unstable gradients; interestingly, this instability does not occur when using our tokenized trajectory reference model.
A.2 Training Details of SMART and CatK.
@@
A.3 Waymo Sim Agent Challenge Metrics
@@
In our experiments, we restrict evaluation to vehicles only: agents corresponding to pedestrians or cyclists are excluded as targets and fixed to their ground-truth trajectories. All reported results are computed on a 2% validation subset of the dataset Zhang et al. (2025a).
A.4 WOSAC Failure Qualitative Results
More qualitative results of WOSAC limitation provided in fig. 4
Figure 4: Additional qualitative examples of WOSAC limitations. Top row: due to sensor noise, the ground-truth agents collided in the data; consequently, safe simulated behaviors are assigned low collision likelihood. Bottom row: the ground-truth agent collided with the road edge, so remaining on-road is given near-zero off-road likelihood and a low map score. These cases illustrate how WOSAC can penalize safe and realistic behaviors when they diverge from noisy or imperfect logged trajectories.
A.5 Planner Variants Overview
@@
A.5.1 Self-Play Policies
@@
A.5.2 Frenet-based Planners
@@
Table 4 summarizes the 10 Frenet-based planner variants, showing their key characteristics including speed ranges, lateral weights, safety focus levels, and sampling densities. These variants range from conservative safety-focused configurations to aggressive high-speed optimized settings.
Table 4: Summary of Frenet-based Planner Variants
@@
Baseline Balanced 0-30 10.0 Medium 15,7,5 Standard configuration
Aggressive High progress 0-35 5.0 Low 15,7,5 Progress weight = 2.0
Conservative Safety-first 0-20 50.0 High 15,7,5 Collision penalty = 5000
Smooth Rider Comfort 0-30 20.0 Medium 15,7,5 Jerk weight = 3.0
Lane Keeper Centerline 0-30 100.0 Medium 15,7,5 Lateral span = 1.5m
@@
Long Horizon Strategic 0-30 10.0 Medium 15,7,5 40 horizon steps
No Collision Test baseline 0-30 10.0 None 15,7,5 Collision disabled
High Speed Highway 5-40 10.0 Medium 15,7,5 Velocity span = 15
@@ A.5.3 IDM-based Planners
Table 5 presents the 10 IDM-based planner variants, each configured to represent different driving styles from cautious urban driving to aggressive highway scenarios. The aggressiveness factor plays a key role in determining the overall behavior of each variant.
Table 5: Summary of IDM-based Planner Variants
@@
Table 6 provides a detailed comparison of the key configuration parameters for representative variants from both planner types, highlighting the differences in their weight distributions and fundamental parameters that lead to their distinct behaviors.
Table 6: Key Configuration Parameters Comparison

### v2

@@ 4.3 SPACeR Agents for Fast & Accurate Closed-Loop Planner Evaluation
A key application of humanlike agents is the closed-loop evaluation of ego-vehicle motion planners. In this setup, the planner under test controls the ego vehicle, while surrounding agents are driven by learned policies. This enables realistic modeling of inter-agent interactions, reactive behaviors, and humanlike decision-making without risking safety.
We evaluate our SPACeR agents along two dimensions: closed-loop evaluation realism and throughput. Unlike distributional realism in section 4.2, which measures similarity to human data, realism here refers to how sim agents change planner rankings under closed-loop interaction.
To assess realism, we compare against two baselines: (i) open-loop log-replay evaluation, and (ii) closed-loop simulation using CAT-K Zhang et al. (2025), a state-of-the-art traffic simulator on WOSAC. We evaluate a diverse set of planners: 18 self-play–trained policies, 10 sampling-based Frenet planners, and 10 IDM-based planners Treiber et al. (2000) (more details can be found in A.7). For each planner, we compute PDM scores Dauner et al. (2024) across multiple scenes under three traffic-simulation strategies: ground-truth logs (GT), CAT-K rollouts, and SPACeR agent policies. We then measure the correlation of PDM scores between strategies. As shown in fig. 4, our approach yields consistently lower correlations with GT replays when compared to CAT-K. We interpret this as evidence that our simulation is more reactive, and suppresses unrealistic planner behavior more effectively—particularly in collision scenarios. Qualitative examples (see fig. 4 and supplementary videos) further highlight that SPACeR agents respond naturally in diverse scenarios, avoiding collisions when reasonable and minimizing unrealistic off-road behaviors.
@@
In summary, SPACeR agents deliver both more realistic closed-loop evaluation and significantly higher throughput, enabling reliable and scalable benchmarking of ego-motion planning policies.
@@ 5 Discussion
Simulating VRUs (Pedestrians and Cyclists). SPACeR improves pedestrian and cyclist realism compared to prior baseline in a large margin, showing that self-play anchoring transfers beyond vehicle agents section 4.2. Nevertheless, VRU self-play remains underexplored: current WOSAC metrics (e.g., collision, off-road) are tailored to vehicles and do not capture sidewalk adherence, crosswalk usage, or other pedestrian specific behaviors. Reward design must likewise reflect these VRU-specific considerations. In this work we primarily use likelihood and collision signals for VRU motion. Developing VRU-aware metrics, reward shaping, and scene-level infrastructure is an important direction for enabling more realistic pedestrian and cyclist simulation.
@@ 6 Conclusion
Acknowledgment
This work was part of W.J. Chang’s summer internship at Applied Intuition, and he is also supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE 2146752. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. The authors would like to thank Eric and Andrew for infrastructure support.
@@ Appendix A Appendix
A.1 Details on SPACeR design choices for simulating VRU
To extend SPACeR to simulating VRUs (pedestrians and cyclists), we ablate several design choices, as shown in table A1. We use separate action heads for each agent type and include a one-hot agent-type indicator in the ego observation so the policy can specialize its behavior. We also find that the reference KL loss consistently improves VRU realism, helping reduce unstable or abrupt pedestrian behaviors. Finally, both the goal-reaching loss and the multi-action head contribute to stable and human-like VRU motion, with noticeable drops in kinematic and interaction realism when either component is removed.
For evaluation, all agents in the scene (including vehicles) are simulated, but WOSAC metrics are computed exclusively for pedestrian and cyclist targets. For training, we follow the same approach as Wu et al. (2024), where different agent types use their own token vocabularies.
Table A1: Ablation study on key components of the SPACeR adapted to VRU simulation settings. Each row removes one design choice. Metrics are reported over VRU target agents only.
Ablation Composite $\uparrow$ Kinematic $\uparrow$ Interactive $\uparrow$ Map-based $\uparrow$ minADE $\downarrow$
Full Model 0.729 0.413 0.762 0.866 2.066
– goal-reaching weight 0.728 0.405 0.769 0.859 2.295
– multi-action head 0.685 0.323 0.742 0.818 3.416
– reference KL loss 0.607 0.222 0.626 0.804 12.844
A.2 Anchoring parameters ablation study
In the ablation study table A2, KL alignment contributes most to realism. While log-likelihood rewards are popular Escontrela et al. (2023), their impact on realism is modest—likely because the real-world driving distribution is highly multi-modal, so maximizing likelihood alone does not yield a stable or sufficiently discriminative signal. A second insight is that, once the policy is anchored to the reference distribution, the explicit goal reward can be removed, which further improves realism.
In fig. A1, we visualize the training dynamics of the anchoring terms. Likelihood-only optimization steadily increases the log-likelihood but also drives entropy downward, indicating reduced diversity as training progresses. In contrast, adding KL alignment not only increases log-likelihood but also keeps the KL term stable and preserves higher entropy, preventing collapse and supporting more reliable top-k realism.
[…]
$r_{\text{task}}$ only 0.70 14.43
Goal + LLH 0.69 21.05
Goal + KL 0.73 4.08
KL + $r_{\text{inf}}$ 0.74 4.73
KL + $r_{\text{inf}}$ + LLH 0.74 4.68
Table A2: Ablations on WOSAC (validation). $r_{\text{inf}}$ = infraction penalties (off-road, collision); LLH = log-likelihood reward; KL alignment essential for realism; goals unnecessary.
Figure A1: Anchoring parameters: effects of likelihood and KL-divergence weights. (a) Likelihood weight effect without KL loss. (b) KL divergence weight sweep without likelihood loss. Likelihood-only optimization increases the probability of the most likely action but collapses entropy, while KL alignment maintains diversity and improves likelihood.
A.3 Training hyperparamters and implementation details of Self-play
The core PPO hyperparameters are summarized in table A3. When training without a reference model, we employ 600 parallel worlds on a single A100 (80 GB, PCIe) GPU. For human-regularized PPO and reference-model training, we reduce the number of worlds to 300, which leads to an approximately $2\times$ slowdown due to the lack of multi-GPU support in GPUDRIVE with madrona backend. With multi-GPU support, we would expect comparable throughput to the reference-free setting. Our policy/value network is configured with an input embedding dimension of 64, a hidden dimension of 128, and a dropout rate of 0.01. To reduce GPU memory consumption in the reference-model setting, we highlight two adjustments: (1) we cap the maximum number of unique scenarios per batch at 200 to increase training speed, and (2) we limit the maximum number of map elements per agent from 200 Kazemkhani et al. (2024) to 120. table A3
Table A3: PPO Training Hyperparameters.
@@ A.3 Training hyperparamters and implementation details of Self-play
We follow the setup of human-regularized PPO Cornelisse and Vinitsky (2024), but adapt it to our tokenized trajectory action space instead of low-level control actions. To train the behavior cloning (BC) reference policy, we use observation–action pairs from the full Waymo Open Motion Dataset (WOMD) rather than the 200-scenario subset originally used, ensuring a stronger expert baseline. Our BC model is parameterized with roughly $2\times$ the capacity of the self-play policy network to increase expressiveness, achieving a validation accuracy of 92%. We implement BC training using the imitation package Gleave et al. (2022) and train for 60 epochs on the full dataset.
During HR-PPO training, we regularize the learned policy against the BC reference policy using a KL-divergence penalty with weight $\beta=0.01$. Larger values of $\beta$ destabilize training, while using tokenized model can generally increase to $\beta=1.0$ w/ similar performance. In addition, when applying human-regularized PPO in the original action space, we found it necessary to clamp the minimum reference log-probability to $10^{-20}$ to avoid unstable gradients; interestingly, this instability does not occur when using our tokenized trajectory reference model.
In the first submission, HR-PPO used the reverse KL divergence, the same as SPACeR. In the camera-ready version, we use the forward KL divergence and report the better-performing variant, without affecting the main conclusions.
A.4 Training Details of SMART and CatK.
@@
A.5 Waymo Sim Agent Challenge Metrics
@@
In our experiments, we restrict evaluation to vehicles only: agents corresponding to pedestrians or cyclists are excluded as targets and fixed to their ground-truth trajectories. All reported results are computed on a 2% validation subset of the dataset Zhang et al. (2025).
A.6 WOSAC Failure Qualitative Results
More qualitative results of WOSAC limitation provided in fig. A2
Figure A2: Additional qualitative examples of WOSAC limitations. Top row: due to sensor noise, the ground-truth agents collided in the data; consequently, safe simulated behaviors are assigned low collision likelihood. Bottom row: the ground-truth agent collided with the road edge, so remaining on-road is given near-zero off-road likelihood and a low map score. These cases illustrate how WOSAC can penalize safe and realistic behaviors when they diverge from noisy or imperfect logged trajectories.
A.7 Planner Variants Overview
@@
A.7.1 Self-Play Policies
@@
A.7.2 Frenet-based Planners
@@
Table A4 summarizes the 10 Frenet-based planner variants, showing their key characteristics including speed ranges, lateral weights, safety focus levels, and sampling densities. These variants range from conservative safety-focused configurations to aggressive high-speed optimized settings.
Table A4: Summary of Frenet-based Planner Variants
@@
Baseline Balanced 0-30 10.0 Medium 10,5,3 Standard configuration
Aggressive High progress 0-35 5.0 Low 10,5,3 Progress weight = 2.0
Conservative Safety-first 0-20 50.0 High 10,5,3 Collision penalty = 5000
Smooth Rider Comfort 0-30 20.0 Medium 10,5,3 Jerk weight = 3.0
Lane Keeper Centerline 0-30 100.0 Medium 15,5,3 Lateral span = 1.5m
@@
Long Horizon Strategic 0-30 10.0 Medium 10,5,3 40 horizon steps
No Collision Test baseline 0-30 10.0 None 10,5,3 Collision disabled
High Speed Highway 5-40 10.0 Medium 10,5,3 Velocity span = 15
@@ A.7.3 IDM-based Planners
Table A5 presents the 10 IDM-based planner variants, each configured to represent different driving styles from cautious urban driving to aggressive highway scenarios. The aggressiveness factor plays a key role in determining the overall behavior of each variant.
Table A5: Summary of IDM-based Planner Variants
@@
Table A6 provides a detailed comparison of the key configuration parameters for representative variants from both planner types, highlighting the differences in their weight distributions and fundamental parameters that lead to their distinct behaviors.
Table A6: Key Configuration Parameters Comparison